OpenStudy (anonymous):

What is the factored form of x^2 + 6x + 8? I got (x+4) (x+2)

OpenStudy (zehanz):

You're right! Just expand again (FOIL)...

OpenStudy (anonymous):

so 5 and 3?

OpenStudy (zehanz):

Or use: 4+2 = 6 (coefficient of x) 4*2 = 8 (constant at the end)

OpenStudy (zehanz):

Rule is: the numbers you are looking for have sum 6 and product 8

OpenStudy (anonymous):

ah thanks so it is 4 and 2 because 4 + 2 = 6 and 4 x 2 = 8

OpenStudy (anonymous):

factoring makes so much sense now thanks

OpenStudy (zehanz):

Yes! Another one: x²+5x+6 = (x+2)(x+3) They're not all that easy though, because sometimes there are negative numbers involved, or there seem to be many numbers to choose from...

OpenStudy (anonymous):

yeah I see it now :)

OpenStudy (zehanz):

E.g. this one: x²-3x-28 Sum = -3 (many possibilities, don't bother with them!) Prod. = -28...hmm that could be: (one of them should be negative, because product is negative ) -1, 28 or 1,-28 no: sum goes wrong -2, 14 or 2, -14 no .... -4, 7 or 4, -7 ...yes! Sum of 4 and -7 is -3 So: (x-7)(x+4) Now I will stop ;)